# BIOPTIC - a target-agnostic efficacy-based small molecules search engine

### v1

@@ Abstract
Recent successes in virtual screening have been made possible by large models and extensive chemical libraries. However, combining these elements is challenging: the larger the model, the more expensive it is to run, making ultra-large libraries unfeasible. To address this, we developed a target-agnostic, efficacy-based molecule search model, which allows us to find structurally dissimilar molecules with similar biological activities. We used the best practices to design fast retrieval system, based on processor-optimized SIMD instructions, enabling us to screen the ultra-large 40B Enamine REAL library with 100% recall rate. We extensively benchmarked our model and several state-of-the-art models for both speed performance and retrieval quality of novel molecules.
@@ 1 Introduction
Combining powerful data-greedy algorithms and ultra-large screening libraries pose a significant challenge. It is realistic to apply fast filters (like Lipinski’s Rule of 5[8], Rule of CNS drugs[9], etc) to huge libraries or to use neural networks to screen small datasets, but screening huge libraries using neural networks takes too much time and money. ML-based drug discovery is a blooming field with a large number of publications. Most of the publications focus on successful case studies rather than how to scale model predictions to huge libraries.
@@ 3 Our contribution
  • We developed a single global model for target-agnostic efficacy-based molecule search, which allows finding structurally dissimilar molecules with similar activities.
@@ 4.1 Modeling
We use binary cross-entropy loss with masking as an objective for learning efficacy-based molecular representations:
@@
for a batch of size $B$, where $\hat{y}_{ij}$ is the prediction of the corresponding molecule $i$ being active to target $j$, and $K$ is the total number of targets used in training. $m_{ij}$ represents a masking indicator, which doesn’t penalize the model on unlabeled data pairs, allowing it to learn from labeled drug-target interactions only.
@@ 4.2 Optimization
After pre-training, the hidden states of the last layer are extracted from the CLS token, and the rest of the tokens are max and average pooled. Three independent linear layers with an output size of 20 are then applied. The reduced representations are concatenated into a vector of size 60. The embedding dimensionality is set to 60 as a trade-off between quality and storage requirements. Embeddings are L2-normalized and passed to the classification linear layer, where the number of classes depends on the target-specific data split and averages around 6700 classes. For demonstrating performance with Hi splits, BIOPTIC is optimized on each data split for a binary classification objective with the LARS optimizer [24], using parameters: $lr=0.1$, weight decay of $31e-$ 3, $\epsilon=1e-8$ and a trust coefficient of $31e-$ 3. To enlarge our dataset, we utilize a randomization of the SMILES strings and random masking of 15% of tokens in 30% of cases during BIOPTIC training.
@@ 5.3 Hi-benchmark results
BIOPTIC 61.4 59.2 61.0 58.2
Table 1: Benchmark results of the models. The best result is in bold. See the metrics description in the Appendix D.

### v2

# BIOPTIC - a target-agnostic potency-based small molecules search engine

@@ Abstract
Recent successes in virtual screening have been made possible by large models and extensive chemical libraries. However, combining these elements is challenging: the larger the model, the more expensive it is to run, making ultra-large libraries unfeasible. To address this, we developed a target-agnostic, potency-based molecule search model, which allows us to find structurally dissimilar molecules with similar biological activities. We used the best practices to design fast retrieval system, based on processor-optimized SIMD instructions, enabling us to screen the ultra-large 40B Enamine REAL library with 100% recall rate. We extensively benchmarked our model and several state-of-the-art models for both speed performance and retrieval quality of novel molecules.
@@ 1 Introduction
Combining powerful data-greedy algorithms and ultra-large screening libraries poses a significant challenge. It is realistic to apply fast filters (like Lipinski’s Rule of 5[8], Rule of CNS drugs[9], etc) to huge libraries or to use neural networks to screen small datasets, but screening huge libraries using neural networks takes too much time and money. ML-based drug discovery is a blooming field with a large number of publications. Most of the publications focus on successful case studies rather than how to scale model predictions to huge libraries.
@@ 3 Our contribution
  • We developed a single global model for target-agnostic potency-based molecule search, which allows finding structurally dissimilar molecules with similar activities.
@@ 4.1 Modeling
We use binary cross-entropy loss with masking as an objective for learning potency-based molecular representations:
@@
for a batch of size $B$, where $\hat{y}_{ij}$ is the prediction of the corresponding molecule $i$ being active to target $j$ (calculated from the logits of the classification layer after applying sigmoid function), $y_{ij}$ is the corresponding ground truth value (0 or 1), and $K$ is the total number of targets used in training. $m_{ij}$ represents a masking indicator, which doesn’t penalize the model on unlabeled data pairs, allowing it to learn from labeled drug-target interactions only.
@@ 4.2 Optimization
After pre-training, the hidden states of the last layer are extracted from the CLS token, and the rest of the tokens are max and average pooled. Three independent linear layers with an output size of 20 are then applied. The reduced representations are concatenated into a vector of size 60. The embedding dimensionality is set to 60 as a trade-off between quality and storage requirements. Embeddings are L2-normalized and passed to the classification linear layer, where the number of classes depends on the target-specific data split and averages around 6700 classes. For demonstrating performance with Hi splits, BIOPTIC is optimized on each data split for a binary classification objective for 300 epochs with the LARS optimizer [24], using parameters: $lr=0.1$, weight decay of $31e-$ 3, $\epsilon=1e-8$ and a trust coefficient of $31e-$ 3. To enlarge our dataset, we utilize a randomization of the SMILES strings and random masking of 15% of tokens in 30% of cases during BIOPTIC training.
@@ 5.3 Hi-benchmark results
BIOPTIC (last) 61.4 59.2 61.0 58.2
BIOPTIC (best) 63.6 61.5 68.0 59.7
Table 1: Benchmark results of the models. Here, checkpoints for Chemprop and Gradient Boosting models are picked at the best Average Precision metric on validation while BIOPTIC was tested in two scenarios: last checkpoint at the end of the training and best checkpoint according to Average Precision on validation. The best result is in bold. See the metrics description in the Appendix D.
@@ 5.5 Retrieval performance on ultra-large molecular libraries
5.6 Scaffold hopping from the box
To demonstrate diverse candidates selection performed with BIOPTIC, we chose ACHE, AA2AR and EGFR as targets and ran search in Enamine REAL Space 40B library. We used 10 diverse active compounds as queries and retrieved top-10k candidates (selection lists). We also took 10k random compounds from Enamine REAL Space and all known molecules tested for binding to the corresponding targets from BindingDB. We then mapped Morgan fingerprints of all the molecules into 2D coordinates with t-SNE. From each selection list, we then removed all molecules that are closer than 0.4 Tanimoto similarity to any BindingDB molecule and picked 10 diverse molecules with RDKit’s MaxMinPicker. The results are shown in 5. It’s seen from the figures that BIOPTIC 1) prioritizes molecules from a widely span corner of the Enamine chemical space, 2) picks molecules structurally different from what is already known, 3) selects candidates of different scaffolds.
Figure 5: Diversity of top candidates retrieved with BIOPTIC from Enamine REAL Space 40 billion library for a) ACHE, b) AA2AR and c) EGFR.